From a building of height h = 16.5 m, you throw a ball of mass of m = 335g downwards with an initial speed vi = 5.7 m/s. What is the final speed vf that the ball is traveling when it hits the ground?
1
Expert's answer
2019-06-07T11:44:20-0400

From the conservation of energy:


"0.5mv_f^2=0.5mv_i^2+mgh"

"v_f^2=v_i^2+2gh=5.7^2+2(9.8)(16.5)"

"v_f=18.9\frac{m}{s}"